OpenStudy (anonymous):

Simplify the given expression. (−3 + 2i) • (2 + i)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

FOIL it out and tell me what you get

OpenStudy (anonymous):

ok brb

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

I got -3+4i is that correct?

jimthompson5910 (jim_thompson5910):

it might help to break (−3 + 2i) • (2 + i) up into -3(2 + i) + 2i(2 + i)

jimthompson5910 (jim_thompson5910):

then you distribute to get -3(2 + i) + 2i(2 + i) -6 -3i + 4i + 2i^2 I'll let you finish up

OpenStudy (anonymous):

-1+2i ! :D

jimthompson5910 (jim_thompson5910):

i^2 = -1, so 2i^2 = 2(-1) = -2

jimthompson5910 (jim_thompson5910):

-6 -3i + 4i + 2i^2 -6 -3i + 4i - 2 what's the last step?

OpenStudy (anonymous):

to add?

jimthompson5910 (jim_thompson5910):

yes to combine like terms

OpenStudy (anonymous):

-8+i? originally I got 7 but thats not an answer choice

jimthompson5910 (jim_thompson5910):

-8+i is correct

OpenStudy (anonymous):

yes ! thankyouu !

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Simplify the given expression. (4 + 2i) − (1 − 7i) im doing this one nowim going to do it and see if i get it right plz

jimthompson5910 (jim_thompson5910):

ok don't forget to distribute fully

OpenStudy (anonymous):

3-5i ?

jimthompson5910 (jim_thompson5910):

(4 + 2i) − (1 − 7i) 4 + 2i − 1 + 7i ... you forgot to distribute the negative fully through now what do you get?

OpenStudy (anonymous):

5+9i !!?

jimthompson5910 (jim_thompson5910):

4-1 is not 5 though

jimthompson5910 (jim_thompson5910):

4 + 2i − 1 + 7i (4-1) + (2i + 7i) ????

OpenStudy (anonymous):

3+9i

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

GOT IT

OpenStudy (anonymous):

@jim_thompson5910 New question ... its same question but its 2 over 2+5i

jimthompson5910 (jim_thompson5910):

the conjugate of 2+5i is 2-5i

jimthompson5910 (jim_thompson5910):

so multiply top and bottom by 2-5i

OpenStudy (anonymous):

I ended up with 4 over 29

jimthompson5910 (jim_thompson5910):

that's incorrect

OpenStudy (anonymous):

and then 2 over 49 those are both choices but idk which one is correct

OpenStudy (anonymous):

the other answers make no sense to me

jimthompson5910 (jim_thompson5910):

what is (2+5i)(2-5i) equal to?

OpenStudy (anonymous):

and when i did the problem came nowhere near

OpenStudy (anonymous):

4 over 10?

OpenStudy (anonymous):

4+10 over 29 ?

jimthompson5910 (jim_thompson5910):

(2+5i)(2-5i) = 2(2-5i) + 5i(2-5i) = ???

OpenStudy (anonymous):

Then its minus 10

OpenStudy (anonymous):

so it would be 4-10i over 29

jimthompson5910 (jim_thompson5910):

4-10i over 29 is correct